# Thermodynamic topology of Black Holes in $F(R)$-Euler-Heisenberg gravity’s Rainbow

### v1

@@ Abstract
The topology of black hole thermodynamics is a fascinating area of study that explores the connections between thermodynamic properties and topological features of black holes. It often involves analyzing critical points in the phase diagrams of black holes and assigning topological charges to these points. One significant approach is based on Duan’s topological current $\phi$-mapping theory, which introduces the concept of topological charges to critical points in black hole thermodynamics. This paper has led to several significant findings: We successfully derive the field equations for $F(R)$-Euler-Heisenberg theory, providing a framework for studying the interplay between modified gravity and non-linear electromagnetic effects. We obtain an analytical solution for a static, spherically symmetric, energy-dependent black hole with constant scalar curvature. Also, our analysis of black holes in F(R)-Euler-Heisenberg gravity’s Rainbow reveals significant insights into their topological properties. We identified the total topological charges by examining the normalized field lines along various free parameters. Our findings indicate that the parameters $(R_{0})$ and $(f_{\epsilon}=g_{\epsilon})$ influence the topological charges. These results are comprehensively summarized in Table I. Additionally, a general overview of Tables II, III, and IV related to the photon sphere of the mentioned black hole reveals that with an increase in $f_{\varepsilon}$, the permissible range of negative $\lambda$ in the first case gradually transitions into a non-permissible region in the third case. On the other hand, it is known that the QED parameter, which measures the strength of nonlinear effects, can be either positive or negative. A positive QED parameter reduces the electric field near the horizon and increases the black hole’s mass, whereas a negative QED parameter increases the electric field and decreases the mass. According to the two statements above, it can be concluded that the increase in $f_{\varepsilon}$ actually decreases the strength of the electric field near the horizon and strengthens the effects of gravity.
@@ I Introduction
The greatest cosmological challenges this century are early-time inflation and late-time acceleration. In precise terms, what does inflation refer to, and what does dark energy consist of? Why and how did the early universe and the late universe expand upwards so very fast? What might explain why the evolution of the universe is similar for small and large curvatures? To tackle these matters, Refs. [1, 2, 3, 4, 5] outline and investigate Einstein’s modified theories of gravity as an alternative framework. Among the various modified theories of gravity, the $F(R)$ theory has aroused a great deal of interest because it offers an explanation for a broad series of observations. For instance, the $F(R)$ theory of gravity provides a plausible hypothesis for the existence of dark matte [6, 7] and the universe’s rapid expansion [2, 3, 4, 5, 8, 9]. Even more, the $F(R)$ theory can be harnessed to forecast cosmic acceleration or early universe inflation [10, 11, 12], not to mention massive compact objects [13, 14, 15, 16, 17, 18]. Another achievement in the context of $F(R)$ gravity is the complete description of all the evolutionary epochs of the universe while maintaining consistency with Newtonian and post-Newtonian approximations [19]. Within the theory of $F(R)$ gravity, $F(R)$ turns out to be an arbitrary function of the scalar curvature $R$.
@@
Since the theory of $F(R)$ gravity may describe a number of cosmological and astrophysical phenomena, it is vital to examine and highlight the coupling of the Euler-Heisenberg (EH) term as a matter source with $F(R)$ gravity’s rainbow by adopting the $F(R)$-EH Lagrangian. Heisenberg and Euler established the non-linear electrodynamics Lagrangian making use of Dirac’s electron-positron theory [20]. Schwinger remodelled this nonperturbative one-loop Lagrangian in the context of quantum electrodynamics (QED) [21] and found that it could identify the phenomenon of vacuum polarisation successfully. Due to the one-loop nonperturbative QED, the EH Lagrangian has drawn ever-increasing scrutiny to the subject of generalised black hole solutions. A wealth of intriguing studies on black holes in the context of EH nonlinear electrodynamics are described in [22, 23, 24, 25, 26, 27, 28, 29].
@@
In the search for the unification of general relativity and quantum mechanics, a wide range of challenges have arisen. Nowadays, in spite of a few theories such as string theory [30], loop quantum gravity [36], and space-time foam models [32] that have attempted to tackle this conundrum, the question remains devoid of any phenomenological process. A key reason for unifying gravity with quantum theory is the violation of Lorentz invariance, which we might regard as an indispensable criterion for the elaboration of a quantum theory of gravity. Another distinctive factor among unification theories is that they forecast a maximum energy value of the order at the Planck scale that a particle could attain; i.e., if a particle is qualified by the Standard Model of particle physics, then it should conform to the upper limit of the Planck energy, $5.10^{19}\textit{e}V$, which is experimentally backed up by Abraham et al [33]. A further alternative to Lorentz symmetry violation involves, in the ultraviolet limit, modifying the standard energy-moment dispersion ratio. This, in fact, has a long history in theories as diverse as discreteness space-time [39], Horava-Lifshitz [34, 35], doubly special relativity [40, 41], and loop quantum gravity [36, 37]. In the framework of doubly special relativity, the dispersion relation works out as $E^{2}f_{\epsilon}^{2}-p^{2}g_{\epsilon}^{2}=m^{2}$ where $f_{\epsilon}$ and $g_{\epsilon}$ are functions of $\epsilon=E/E_{P}$, $E$ standing for the energy of the particle used to analyse space-time and $E_{P}$ for Planck’s energy. The functions $f_{\epsilon}$ and $g_{\epsilon}$ are commonly thought of as rainbow functions and are designed for phenomenological purposes [42]. On the other hand, the standard energy dispersion relation gets recurved in the limit $\lim\limits_{\epsilon\to 0}f_{\epsilon}=\lim\limits_{\epsilon\to 0}g_{\epsilon}=1$, otherwise known as the infrared limit. As a matter of fact, Linear Lorentz transformations break the invariance of the theory [43]; nevertheless, non-linear Lorentz transformations are those that keep the theory invariant. Moreover, as well as the speed of light being a constant, the Planck energy is also a constant, and it remains impossible for a particle to rise beyond this limit.
@@
A generalisation of doubly special relativity to what is known as rainbow gravity has been suggested by Magueijo and Smolin [44], where spacetime is mapped to a family of specified parameters in the metric that is parametrised by $\epsilon$. Accordingly, the spacetime geometry is contingent on the energy of the particle testing it, causing a rainbow of the metric. Meanwhile, a number of studies have incorporated Magueijo and Smolin’s theory into the understanding of black holes [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] and into modifying theories of gravity [60, 61]. A worthy concrete illustration investigation is concerned with the work of M. Momennia et al. [62] exploiting a non-linear electrodynamics (NED) field in the context of the rainbow gravity, resulting in the validity of the first law of thermodynamics in the presence of rainbow functions. Therefore, the rainbow gravity framework has provided the avenue for the consideration of a large number of investigations, such as the derivation considering the dilatonic gravity of an exact black hole solution minimally coupled to the Born-Infeld term with an energy-dependent Liouville-type potential [63]. Similarly, incorporating the quadratic term of the Gauss-Bonnet gravity in the framework of rainbow gravity leads to modifying the appropriate constraints to work out non-singular universes [64]. It is worth noting that the study of thermodynamic properties is considered the main and impacting process regarding rainbow gravity. Thus, since the energy scale depends on the spacetime metric, entropy and temperature also have the same dependence [65]. To gain a topological perspective in thermodynamics, one effective approach is to use Duans topological current $\phi$ mapping theory. Wei et al. introduced two distinct methods to study topological thermodynamics based on temperature and the generalized free energy function. The first method involves analyzing the temperature function by eliminating pressure and utilizing the auxiliary and topological parameter $1/\sin\theta$. The potential is then constructed based on these assumptions. For further study, see [66, 67, 68, 69]. The second method assumes that black holes can be considered defects in the thermodynamic parameter space. Their solutions are investigated using the generalized off-shell free energy. In this context, the stability and instability of the obtained black hole solutions are determined by positive and negative winding numbers, respectively. Additionally, the properties of a field configuration can be deduced from the zero points of the field in space [70, 71, 72, 73]. Wei, Liu, and B. Mann asserted that for different branches of a black hole at an arbitrary temperature, the topological charge number, the sum of the winding numbers, is a universal number independent of the black hole parameters. In the range of small and large black holes, the topological charge number depends solely on the thermodynamic asymptotic behavior of the black hole temperature. Consequently, black holes can be categorized into three different groups based on the number of topological charges [70]. Following the introduction of this second method, various black holes were studied to investigate topological charges. C. H. Liu and J. Wang demonstrated that the topological number remains constant, and the charged GB black holes in the AdS space fall into the same category as the RN-AdS, sharing the same topological numbers. However, their results indicated that charge dependence is not the only effective factor [74]. Y. Du and Xi. Zhang examined the rotating charged BTZ black hole model and found that there are only two topological classes for BTZ space-time. [75]. Using the same method, Wu investigated the structure of charged Lorentzian Taub-NUT space-times and neutral Lorentzian NUT-charged space-times in 4 dimensions in two separate articles. [76]. In the second work, Wu found that the presence of the NUT charge as a parameter did not affect the structures in symptotically flat space-time but had an effect on asymptotically local AdS space-time [77]. For further study, you can see [78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95]. Building upon the above concepts, we organize the article as follows: In Section II, we provide the field equations in F(R)-Euler-Heisenberg theory. In Section III, we obtain the black hole solutions in F(R)-Euler-Heisenberg gravity’s rainbow. In Sections IV and V, we study the thermodynamics, thermodynamic topology, and photon sphere of the mentioned black holes. We obtain the thermodynamic topology and topological classification of black holes in F(R)-Euler-Heisenberg gravity’s rainbow and its photon sphere in some cases in Section VI. Finally, we present the conclusion of our paper in Section VII.
@@ II Field equations in $F(R)$-Euler-Heisenberg theory
where $F(R)=R+f(R)$, in which $R$ and $f(R)$, are respectively, the Ricci scalar and a general function of the Ricci scalar, $g=\det(g_{\mu\nu})$ is the determinant of the metric tensor $g_{\mu\nu}$, and $\mathcal{L}(\mathcal{S},\mathcal{Q})$ is dedicated to being the EH Lagrangian. Throughout this study, we adopt the Newtonian gravitational constant and the speed of light as being equal to 1, i.e., $G=c=1$. In practical terms, the EH Lagrangian can be described by considering the following expression [20]:
@@ III Black hole solutions in $F(R)$-Euler-Heisenberg gravity’s rainbow
Within the scope of this study, we consider a spherically, static, four-dimensional, energy-dependent spacetime, in which we adopt the procedure described in Refs. [44, 101], so that
@@
Practically speaking, the field equations underlying the context of $F(R)$ gravity with the EH matter fields (9) seem to be difficult. This means that accurate analytical solutions are proving tricky to pin down. To overcome this issue, one should consider the traceless energy-momentum tensor for the EH matter field, and then one can obtain an analytical solution from $F(R)$ gravity coupled to a nonlinear electrodynamics field. Thus, to be able to derive the solution of a black hole with constant curvature in the $F(R)$ theory of gravity as coupled to the EH matter field, it suffices for the trace of the stress-energy tensor $T_{\mu\nu}$ to hold to zero [102, 103]. Based on the hypothesis of constant scalar curvature $R=R_{0}=\text{constant}$ [104], the trace of equation (9) takes the following terms:
@@
By way of observation, the presence of the spacetime singularity is due to the mass $(m_{0})$, the electric charge $(q)$, and the EH parameter $(\lambda)$ in the black hole metric. A useful way of avoiding the singularity resulting from the mass, the charge, and the EH parameters is to outline a procedure with a non-linear charge distribution function similar to Ref. [105]. Throughout this work, we will not be thinking about such a situation and will stick to the metric function (14) for the rest of the analysis. On the other hand, examining behavior at large distances is a worthwhile adjunct, given that

### v2

@@ Abstract
The topology of black hole thermodynamics is a fascinating area of study that explores the connections between thermodynamic properties and topological features of black holes. It often involves analyzing critical points in the phase diagrams of black holes and assigning topological charges to these points. One significant approach is based on Duan’s topological current $\phi$-mapping theory, which introduces the concept of topological charges to critical points in black hole thermodynamics. We successfully derive the field equations for $F(R)$-Euler-Heisenberg theory, providing a framework for studying the interplay between modified gravity and non-linear electromagnetic effects. We obtain an analytical solution for a static, spherically symmetric, energy-dependent black hole with constant scalar curvature. Also, our analysis of black holes in F(R)-Euler-Heisenberg gravity’s Rainbow reveals significant insights into their topological properties. We identified the total topological charges by examining the normalized field lines along various free parameters. Our findings indicate that the parameters $(R_{0})$ and $(f_{\epsilon}=g_{\epsilon})$ influence the topological charges. These results are comprehensively summarized in Table I. In examining the photon sphere within this model, the sign of the parameter $R_{0}$ plays a crucial role in determining whether the model adopts a dS or AdS configuration. An interesting characteristic of this model is that, in its AdS form, it avoids the formation of naked singularity regions, which sets it apart from many other models. Typically, varying parameter values in other models can result in the division of space into regions of black holes and naked singularities. However, this model consistently retains its black hole behavior by featuring an unstable photon sphere, regardless of parameter values within the acceptable range. In its dS form, the behavior of the model’s photon sphere remains consistent with other dS models and does not exhibit unique differences.
Thermodynamic topology; F(R)-Euler-Heisenberg gravity’s Rainbow, Photon sphere
@@ I Introduction
The greatest cosmological challenges this century are early-time inflation and late-time acceleration. In precise terms, what does inflation refer to, and what does dark energy consist of? Why and how did the early universe and the late universe expand upwards so very fast? What might explain why the evolution of the universe is similar for small and large curvatures? To tackle these matters, Refs. DeFelice:2010aj ; Faraoni:2010pgm ; Nojiri:2010wj ; Capozziello:2011et ; Bamba:2012cp outline and investigate Einstein’s modified theories of gravity as an alternative framework. Among the various modified theories of gravity, the $F(R)$ theory has aroused a great deal of interest because it offers an explanation for a broad series of observations. For instance, the $F(R)$ theory of gravity provides a plausible hypothesis for the existence of dark matte Abdalla:2020ypg ; Nojiri:2006ri and the universe’s rapid expansion Faraoni:2010pgm ; Nojiri:2010wj ; Capozziello:2011et ; Bamba:2012cp ; Joyce:2014kja ; Nojiri:2017ncd . Even more, the $F(R)$ theory can be harnessed to forecast cosmic acceleration or early universe inflation Capozziello:2002rd ; Aditya:2018cmn ; Hu:2007nk , not to mention massive compact objects Cooney:2009rr ; Cheoun:2013tsa ; Astashenok:2020cfv ; Astashenok:2021peo ; Sarmah:2021ule ; Kalita:2021zjg . Another achievement in the context of $F(R)$ gravity is the complete description of all the evolutionary epochs of the universe while maintaining consistency with Newtonian and post-Newtonian approximations Capozziello:2007ms . Within the theory of $F(R)$ gravity, $F(R)$ turns out to be an arbitrary function of the scalar curvature $R$.
@@
Since the theory of $F(R)$ gravity may describe a number of cosmological and astrophysical phenomena, it is vital to examine and highlight the coupling of the Euler-Heisenberg (EH) term as a matter source with $F(R)$ gravity’s rainbow by adopting the $F(R)$-EH Lagrangian. Heisenberg and Euler established the non-linear electrodynamics Lagrangian making use of Dirac’s electron-positron theory Heisenberg:1936nmg . Schwinger remodelled this nonperturbative one-loop Lagrangian in the context of quantum electrodynamics (QED) Schwinger:1951nm and found that it could identify the phenomenon of vacuum polarisation successfully. Due to the one-loop nonperturbative QED, the EH Lagrangian has drawn ever-increasing scrutiny to the subject of generalised black hole solutions. A wealth of intriguing studies on black holes in the context of EH nonlinear electrodynamics are described in Bronnikov:1979ex ; Bronnikov:2000vy ; Yajima:2000kw ; Corda:2009xd ; Ruffini:2013hia ; Hendi:2014xia ; Maceda:2018zim ; Olvera:2019unw .
@@
In the search for the unification of general relativity and quantum mechanics, a wide range of challenges have arisen. Nowadays, in spite of a few theories such as string theory Amati:1988tn , loop quantum gravity Amelino-Camelia:1996bln , and space-time foam models Amelino-Camelia:1997ieq that have attempted to tackle this conundrum, the question remains devoid of any phenomenological process. A key reason for unifying gravity with quantum theory is the violation of Lorentz invariance, which we might regard as an indispensable criterion for the elaboration of a quantum theory of gravity. Another distinctive factor among unification theories is that they forecast a maximum energy value of the order at the Planck scale that a particle could attain; i.e., if a particle is qualified by the Standard Model of particle physics, then it should conform to the upper limit of the Planck energy, $5.10^{19}\textit{e}V$, which is experimentally backed up by Abraham et al PierreAuger:2010gfm . A further alternative to Lorentz symmetry violation involves, in the ultraviolet limit, modifying the standard energy-moment dispersion ratio. This, in fact, has a long history in theories as diverse as discreteness space-time tHooft:1996ziz , Horava-Lifshitz Horava:2009uw ; Horava:2009if , doubly special relativity Magueijo:2002am ; Magueijo:2001cr , and loop quantum gravity Amelino-Camelia:1996bln ; Amelino-Camelia:2008aez . In the framework of doubly special relativity, the dispersion relation works out as $E^{2}f_{\epsilon}^{2}-p^{2}g_{\epsilon}^{2}=m^{2}$ where $f_{\epsilon}$ and $g_{\epsilon}$ are functions of $\epsilon=E/E_{P}$, $E$ standing for the energy of the particle used to analyse space-time and $E_{P}$ for Planck’s energy. The functions $f_{\epsilon}$ and $g_{\epsilon}$ are commonly thought of as rainbow functions and are designed for phenomenological purposes Dehghani:2018svw . On the other hand, the standard energy dispersion relation gets recurved in the limit $\lim\limits_{\epsilon\to 0}f_{\epsilon}=\lim\limits_{\epsilon\to 0}g_{\epsilon}=1$, otherwise known as the infrared limit. As a matter of fact, Linear Lorentz transformations break the invariance of the theory Garattini:2014rwa ; nevertheless, non-linear Lorentz transformations are those that keep the theory invariant. Moreover, as well as the speed of light being a constant, the Planck energy is also a constant, and it remains impossible for a particle to rise beyond this limit.
@@
A generalisation of doubly special relativity to what is known as rainbow gravity has been suggested by Magueijo and Smolin Magueijo:2002xx , where spacetime is mapped to a family of specified parameters in the metric that is parametrised by $\epsilon$. Accordingly, the spacetime geometry is contingent on the energy of the particle testing it, causing a rainbow of the metric. Meanwhile, a number of studies have incorporated Magueijo and Smolin’s theory into the understanding of black holes a1 ; a2 ; a3 ; a4 ; a5 ; a6 ; a7 ; a8 ; a9 ; a10 ; a11 ; a12 ; a13 ; a14 ; a15 and into modifying theories of gravity Garattini:2012ec ; Hendi:2016oxk . A worthy concrete illustration investigation is concerned with the work of M. Momennia et al. Hendi:2015pwk exploiting a non-linear electrodynamics (NED) field in the context of the rainbow gravity, resulting in the validity of the first law of thermodynamics in the presence of rainbow functions. Therefore, the rainbow gravity framework has provided the avenue for the consideration of a large number of investigations, such as the derivation considering the dilatonic gravity of an exact black hole solution minimally coupled to the Born-Infeld term with an energy-dependent Liouville-type potential Hendi:2017ptl . Similarly, incorporating the quadratic term of the Gauss-Bonnet gravity in the framework of rainbow gravity leads to modifying the appropriate constraints to work out non-singular universes Hendi:2016tiy . It is worth noting that the study of thermodynamic properties is considered the main and impacting process regarding rainbow gravity. Thus, since the energy scale depends on the spacetime metric, entropy and temperature also have the same dependence Dehghani:2018qvn . To gain a topological perspective in thermodynamics, one effective approach is to use Duans topological current $\phi$ mapping theory. Wei et al. introduced two distinct methods to study topological thermodynamics based on temperature and the generalized free energy function. The first method involves analyzing the temperature function by eliminating pressure and utilizing the auxiliary and topological parameter $1/\sin\theta$. The potential is then constructed based on these assumptions. For further study, see sa1 ; sa2 ; sa3 ; sa4 . The second method assumes that black holes can be considered defects in the thermodynamic parameter space. Their solutions are investigated using the generalized off-shell free energy. In this context, the stability and instability of the obtained black hole solutions are determined by positive and negative winding numbers, respectively. Additionally, the properties of a field configuration can be deduced from the zero points of the field in space sa5 ; sa6 ; sa7 ; sa8 . Wei, Liu, and B. Mann asserted that for different branches of a black hole at an arbitrary temperature, the topological charge number, the sum of the winding numbers, is a universal number independent of the black hole parameters. In the range of small and large black holes, the topological charge number depends solely on the thermodynamic asymptotic behavior of the black hole temperature. Consequently, black holes can be categorized into three different groups based on the number of topological charges sa5 . Following the introduction of this second method, various black holes were studied to investigate topological charges. C. H. Liu and J. Wang demonstrated that the topological number remains constant, and the charged GB black holes in the AdS space fall into the same category as the RN-AdS, sharing the same topological numbers. However, their results indicated that charge dependence is not the only effective factor sa9 . Y. Du and Xi. Zhang examined the rotating charged BTZ black hole model and found that there are only two topological classes for BTZ space-time. sa10 . Using the same method, Wu investigated the structure of charged Lorentzian Taub-NUT space-times and neutral Lorentzian NUT-charged space-times in 4 dimensions in two separate articles. sa11 . In the second work, Wu found that the presence of the NUT charge as a parameter did not affect the structures in symptotically flat space-time but had an effect on asymptotically local AdS space-time sa12 . For further study, you can see sa13 ; sa14 ; sa15 ; sa16 ; sa17 ; sa18 ; sa19 ; sa20 ; sa21 ; sa22 ; sa23 ; sa24 ; sa25 ; sa26 ; sa27 ; sa28 ; sa29 ; sa30 ; 500 ; 501 ; 502 ; 503 ; 504 ; 505 ; 506 ; 507 . Building upon the above concepts, we organize the article as follows: In Section II, we provide the field equations in F(R)-Euler-Heisenberg theory. In Section III, we obtain the black hole solutions in F(R)-Euler-Heisenberg gravity’s rainbow. In Sections IV and V, we study the thermodynamics, thermodynamic topology, and photon sphere of the mentioned black holes. We obtain the thermodynamic topology and topological classification of black holes in F(R)-Euler-Heisenberg gravity’s rainbow and its photon sphere in some cases in Section VI. Finally, we present the conclusion of our paper in Section VII.
@@ II Field equations in $F(R)$-Euler-Heisenberg theory
where $F(R)=R+f(R)$, in which $R$ and $f(R)$, are respectively, the Ricci scalar and a general function of the Ricci scalar, $g=\det(g_{\mu\nu})$ is the determinant of the metric tensor $g_{\mu\nu}$, and $\mathcal{L}(\mathcal{S},\mathcal{Q})$ is dedicated to being the EH Lagrangian. Throughout this study, we adopt the Newtonian gravitational constant and the speed of light as being equal to 1, i.e., $G=c=1$. In practical terms, the EH Lagrangian can be described by considering the following expression Heisenberg:1936nmg :
@@ III Black hole solutions in $F(R)$-Euler-Heisenberg gravity’s rainbow
Within the scope of this study, we consider a spherically, static, four-dimensional, energy-dependent spacetime, in which we adopt the procedure described in Refs. Magueijo:2002xx ; peng2008covariant , so that
@@
Practically speaking, the field equations underlying the context of $F(R)$ gravity with the EH matter fields (9) seem to be difficult. This means that accurate analytical solutions are proving tricky to pin down. To overcome this issue, one should consider the traceless energy-momentum tensor for the EH matter field, and then one can obtain an analytical solution from $F(R)$ gravity coupled to a nonlinear electrodynamics field. Thus, to be able to derive the solution of a black hole with constant curvature in the $F(R)$ theory of gravity as coupled to the EH matter field, it suffices for the trace of the stress-energy tensor $T_{\mu\nu}$ to hold to zero delaCruz-Dombriz:2009pzc ; Moon:2011hq . Based on the hypothesis of constant scalar curvature $R=R_{0}=\text{constant}$ Cognola:2005de , the trace of equation (9) takes the following terms:
@@
By way of observation, the presence of the spacetime singularity is due to the mass $(m_{0})$, the electric charge $(q)$, and the EH parameter $(\lambda)$ in the black hole metric. A useful way of avoiding the singularity resulting from the mass, the charge, and the EH parameters is to outline a procedure with a non-linear charge distribution function similar to Ref. Balart:2014cga . Throughout this work, we will not be thinking about such a situation and will stick to the metric function (14) for the rest of the analysis. On the other hand, examining behavior at large distances is a worthwhile adjunct, given that